OpenStudy (jurynee):

Mark, Taylor, & Brian's ages are consecutive whole numbers. Mark is the youngest, and Brian is the oldest. The sum of their ages is 111. How old is Taylor? A) 34 years old B) 35 years old C) 36 years old D) 37 years old

OpenStudy (photon336):

consecutive means one after the other. \[Mark = n\] \[Taylor, n+1\] \[Brian, n+2\] now the SUM = addition. just solve for n \[n+(n+1)+(n+2) = 3n+3 = 111\]

OpenStudy (photon336):

@jurynee so what's n or mark's age?

OpenStudy (jurynee):

im trying to solve it out

OpenStudy (jurynee):

would it be 6

OpenStudy (photon336):

so we've got 3n+3 = 111 we need to get 3n by itself we can do this by subtracting 3 from both sides first

OpenStudy (jurynee):

nvm

OpenStudy (photon336):

\[3n = 108 \] so what's n now marks age?

OpenStudy (jurynee):

36

OpenStudy (photon336):

now what's 36+37+38?

OpenStudy (jurynee):

111

OpenStudy (jurynee):

i got it wrong it was 37

OpenStudy (photon336):

it asked for taylor's age.. mark is 36 so taylor would be 37 and the other guy 38

OpenStudy (photon336):

did you understand how I got the equation?

OpenStudy (jurynee):

yes